# On the virialization threshold for halo mass functions

### v1

@@ 4 Sheth-Tormen halo mass function with $\delta_{{\rm v}}$
Now let us refit these parameters with the substitution of $\delta_{c}$ by $\delta_{{\rm v}}$. First, we mention that a simple way to substitute $\delta_{c}$ by $\delta_{{\rm v}}$ was given [26], where the original parameter $a=0.707$ was rescaled such that the EdS limit is the same, so
@@
Therefore, as can be seen from the values in (4.5), refitting the ST-HMF with $\delta_{{\rm v}}$ produce parameters $a$ and $p$ that are much closer to the PS values than the original values or the more recent fit (4.2), while the parameter $A$ has a very small change. Moreover, even when fitting for larger mass intervals, the parameters $a$ and $p$ still reach values closer to the PS limit when compared to the values of (4.2). This a concrete example that the use of $\delta_{{\rm v}}$ can provide HMF less dependent of external parameters.
@@
In figure 3, we show the percent differences of ST and STv-HMFs against the E-HMF. The general trend is clear: the STv-HMF has better agreement with E-HFM than ST-HMF, expect for some small intermediate mass interval. In particular, at $z=0$ and for $\Omega_{m}^{0}=0.3$, the deviation of STv from E-HMF is smaller than $1\%$ for $M<1.8\times 10^{14}M_{\odot}h^{-1}$ and for $M>4.4\times 10^{14}M_{\odot}h^{-1}$ it is always smaller then ST-HMF.
@@
Figure 3: Differences between ST-HMF with parameters fitted using $\delta_{c}$ (ST, solid lines), equation (4.2), and with $\delta_{{\rm v}}$ (STv, dashed lines), equation (4.5), with respect to the E-HMF, equation (3.3). Each panel shows the differences for distinct values of $\Omega_{m}^{0}$, indicated in the title of each plot, and for three redshifts indicated in the legend of the right panel.
@@
Besides the improved agreement with the E-HMF, it is also important to note that, as expressed in [6], the residuals between their model and their numerical HMF “are larger at high $\nu$, where the best fit predicts more halos than we find in the simulations at $z=0$.” Since $\delta_{{\rm v}}$ increases at low-$z$, whereas $\delta_{c}$ decreases, it is clear that a HMF with the virialization threshold is more suitable. This is also represented by the better agreement between STv and E-HMFs at higher masses, which occurs for all redshifts and values of $\Omega_{m}^{0}$ shown in figure 3.
@@ 5 Beyond $\Lambda$CDM models
As a further consistency check of the use of $\delta_{{\rm v}}$ as the threshold for halo formation in the ST template, we will analyse the level of agreement between E, ST and STv-HMFs when computing the impact of smooth DE models with $w\neq-1$. HMFs for smooth DE with varying equation of state parameter have been studied by [7], where the effects were modelled by an explicit dependence on the matter growth rate, $f=d\ln\delta_{m}/d\ln a$. However, as shown in [8], this HMF has an overall worse agreement with the E-HMF than the ST-HMF. It is also discussed that the dependence on $f$ can be reparameterised in terms of $\Omega_{m}$, which appears in equation (3.5), making use of growth index parametrization $f=\Omega_{m}^{\gamma}$. Therefore, although the E-HMF has been fitted for the $\Lambda$CDM model, one can expect that the impact of $w\neq-1$ can be consistently captured, and it will be used here as a basis of comparison.
@@ 5.1 Constant EoS
Figure 4: $R_{n}$, defined in (5.2), for STv, ST and E-HMFs (indicated in the plot legends) for dark energy models with $w_{0}=-0.8$ (left panel) and $w_{0}=-1.2$ (right panel) at $z=0$ (blue lines) and $z=0.5$ (orange lines). (5.2)
An important feature to note in these examples is the clear impact of phantom ($w<-1$) and non-phaton EoS ($w>-1$) with respect to $\Lambda$CDM. For $w_{0}=-0.8$, the DE fluid has more energy density than $\Lambda$ at high-$z$, consequently $\Omega_{m}\left(z\right)$ is lower and the growth of matter fluctuations is relatively suppressed. As a consequence, we always have $R_{n}<1$. The opposite occurs for the phantom case of $w_{0}=-1.2$.
@@ 5.2 Time varying EoS
Although the impact of constant EoS is very clear, and all the three HMFs agree qualitatively about it, the situation for time-varying EoS can be more complex. As an example, we will analyse two pairs of $\left(w_{0},w_{a}\right)$ values, whose growth history changes in time with respect to the $\Lambda$CDM one. For $p_{1}=\left(w_{0}=-0.9,w_{a}=-0.3\right)$ the EoS is phantom at high-$z$ and non-phatom at low-$z$. For $p_{2}=\left(w_{0}=-1.104,w_{a}=0.3\right)$, the converse holds. It is interesting to note that the first kind of EoS evolution is supported by the recent DESI-BAO analysis [40].
Let us first analyse the relative linear growth history with respect to $\Lambda$CDM. In figure 5, we show the ratio between the evolutions of $\sigma_{8}\left(z\right)$, given by
@@
For $p_{1}$, $\sigma_{8}$ is higher than in $\Lambda$CDM at high-$z$. At lower $z$, the EoS becomes non-phatom and there is less growth with respect to $\Lambda$CDM. The opposite happens for $p_{2}$. In both cases, this relative transition in the growth behaviour occurs at $z_{t}\simeq 0.38$ (the value of $w_{0}$ in $p_{2}$ was adjusted to get approximately the same $z_{t}$ as for $p_{1}$).
Figure 5: Evolution of $R_{\sigma_{8}}\left(z\right)$, defined in (5.4), for two pairs of $w_{0}$ and $w_{a}$ indicated in the plot legend.
Prior to $z_{t}\simeq 0.38$, the growth histories of the two examples are clearly enhanced or diminished with respect to $\Lambda$CDM. It is not expected that the nonlinear evolution encoded in the effective threshold of HMFs should present a distinct relative behaviour as seen for $R_{\sigma_{8}}$. Therefore, at $z>z_{t}$, the ratio between halo number densities, equation (5.2), should preserve the relations seen in figure 5.
However, these relations between models based on the linear growth history is not always satisfied by HMFs. In figure 6 we show $R_{n}$ for the two models considered in this subsection. As can be seen, at $z=0.4$, slightly before $z_{t}\simeq 0.38$, only ST-HMF presents the expected behaviour: $R_{n}>1$ for $p_{1}$ and $R_{n}<1$ for $p_{2}$. This is easy to understand, given that the effective threshold in the ST-HMF depends on $\delta_{c}$, which is nearly constant, the main impact in $R_{n}$ is due to $\sigma\left(M,z\right)$. For STv and E-HMF, the effective thresholds vary in time, but in a manner not totally consistent with the variation of $\sigma_{8}\left(z\right)$.
@@
Figure 6: Ratios of STv, ST and E-HMFs (indicated in the plot legends) for DE models with EoS parameters indicated at the plot titles at $z=0$ (blue lines) and $z=1$ (orange lines).
In order to clearly see that this undesirable effect is caused by the effective threshold, in figure 7, we show the redshift evolution of the ratios
$R_{c}=\frac{D_{c\,\text{mod}}}{D_{c\,\Lambda}}\,\,\,\text{and }R_{{\rm v}}=\frac{\delta_{{\rm v\,}\text{mod}}}{\delta_{{\rm v}\,\Lambda}}\,,$ (5.5)
where $D_{c}$ is the time-dependent part of the effective threshold in the E-HMF, defined in (3.13). At high-$z$, both quantities are lower than unity for $p_{1}$ and higher than unity for $p_{2}$, as expected for enhanced and diminished growth histories with respect to $\Lambda$CDM, respectively. The inversion in the relative behaviour occurs much earlier than the one related to $\sigma_{8}\left(z\right)$, shown in figure 5. After this inversion, $R_{\rm v}$ remains above or lower the unity, whereas $R_{c}\rightarrow 1$ as $z\rightarrow 0$. Moreover, $R_{c}$ reaches higher or lower values than $R_{\rm v}$, with a stronger time variation.
Figure 7: Evolution of $R_{c}-1$ and $R_{{\rm v}}-1$, defined in (5.5), for $p_{1}=\left(w_{0}=-0.9,w_{a}=-0.3\right)$ and $p_{2}=\left(w_{0}=-1.104,w_{a}=0.3\right)$.
@@
Moreover, $R_{c}$ shows a second problem. Given that the redshift evolution of the effective threshold in the E-HMF mainly depends on $\Omega_{m}^{a_{z}}$, for fixed $\Omega_{m}^{0}$, $R_{c}$ for different models have the same value at $z=0$. This looks unnatural and is in disagreement with behaviour $R_{\sigma_{8}}$ shown in figure 5. Such behaviour would indicate that, although a specific model has more/less linear growth than $\Lambda$CDM at low-$z$, the effective threshold does not respond to this increase/decrease of growth at low-$z$. This issue might indicate that the parameter $a_{z}$ in the E-HMF should have some dependence on the EoS. On the other hand, $R_{{\rm v}}$ reproduces the behaviour of $R_{\sigma_{8}}$ at low-$z$, what looks more natural.
@@ 6 Conclusions
We have analysed the redshift dependence of the effective threshold in the E-HMF, proposed in [8], showing it is qualitatively equivalent to the virialization threshold, $\delta_{{\rm v}}$, and incompatible with the evolution of the collapse threshold, $\delta_{c}$, which has been used in many HMFs so far. This finding indicates that the abundance of halos in the Universe is better described by the physics of virialization, encoded in $\delta_{{\rm v}}$, than physics of collapse, associated with $\delta_{c}$.
We found that the redshift-dependent part of the E-HMF, $\Omega_{m}^{a_{z}}\delta_{c}^{2}$, can be accurately represented by a simple power-law function of the virialization threshold, given by $\alpha\delta_{{\rm v}}^{\beta}$, which is also robust against variations of $\Omega_{m}^{0}$, showing that the natural dependence of the threshold on $\delta_{{\rm v}}$ is not a coincidence for a particular expansion history in the $\Lambda$CDM model. We found $\beta\simeq 4.37$, which shows that the effective threshold in the E-HMF grows faster than $\delta_{{\rm v}}^{2}$ at low-$z$. We interpret this as an indication that a more realistic model for the nonlinear dynamics and virialization should provide faster growing $\delta_{{\rm v}}$, in such a way that effective threshold is proportional to the square of this new function, as demanded by the analytical formulation of the HMF.
Exploring these findings, we have shown that a refitted ST-HMF with the substitution $\delta_{c}\rightarrow\delta_{{\rm v}}$ has better agreement with the E-HMF, which is also robust when varying $\Omega_{m}^{0}$. Interestingly, this refit increases the parameter $a$ and decreases the parameter $p$, showing that the deviation of the ST-HMF with $\delta_{{\rm v}}$ from the PS-HMF is smaller. This is a concrete example that the use of the virialization threshold in HMFs is more natural and can diminish their dependence on extra parameters.
@@
Our main conclusion is that the virialization threshold is a sensible basic model for the description of the time-dependence of the effective threshold for smooth DE models. Given that $\delta_{{\rm v}}$ is computed by a physical model, in principle, it should provide a reliable manner to determine the impact models beyond smooth DE in halo abundances, such as clustering DE, or even for the scenarios of modified gravity and models beyond cold dark matter. Furthermore, at least in the case of clustering DE, it is not numerically possible to compute $\delta_{c}$ [29], thus the use of $\delta_{{\rm v}}$ is essential to explore the behaviour of such models with semi-analitic HMFs.

### v2

@@ 4 Sheth-Tormen halo mass function with $\delta_{{\rm v}}$
Now let us refit these parameters substituting $\delta_{c}(z)$ by $\delta_{{\rm v}}(z)$. First, we mention that a simple way to substitute $\delta_{c}$ by $\delta_{{\rm v}}$ was given [26], where the original parameter $a=0.707$ was rescaled such that the EdS limit is the same, so
@@
Therefore, as can be seen from the values in (4.5), refitting the ST-HMF with $\delta_{{\rm v}}$ produces parameters $a$ and $p$ that are much closer to the PS values than the original values or the more recent fit (4.2), while the parameter $A$ has a very small change. Moreover, even when fitting for larger mass intervals, the parameters $a$ and $p$ still reach values closer to the PS limit when compared to the values of (4.2). Since the PS-HMF is an analytical formulation, this a concrete example that the use of $\delta_{{\rm v}}$ as basic threshold function can provide a HMF less dependent on ad doc parameters, while providing better agreement with the E-HMF.
@@
In figure 3, we show the percent differences of ST and STv-HMFs against the E-HMF. The general trend is clear: the STv-HMF has better agreement with E-HFM than ST-HMF, expect for some small intermediate mass interval. In particular, at $z=0$ and for $\Omega_{m0}=0.3$, the deviation of STv from E-HMF is smaller than $1\%$ for $M<1.8\times 10^{14}M_{\odot}h^{-1}$ and for $M>4.4\times 10^{14}M_{\odot}h^{-1}$ it is always smaller then ST-HMF.
@@
Figure 3: Differences between ST template with respect to the E-HMF, equation (3.3), for three sets o parameters. The ST label indicates the parameters given by Despali et al. 2016, ref. [6], (solid lines), equation (4.2). The STv label indicates our refit using $\delta_{{\rm v}}$ (dashed lines), equation (4.5), and STc the refit of the ST template using $\delta_{c}$, done in the same fashion as for the STv case (dotted lines). Each panel shows the differences for distinct values of $\Omega_{m0}$, indicated in the title of each plot, and for three redshifts indicated in the legend of the right panel. The y-scale is linear between ±$\pm 10\%$ and symmetric-log beyond these values. The gray bands indicate ±$\pm 5\%$ region.
@@
Besides the improved agreement with the E-HMF, it is also important to note that, as expressed in [6], the residuals between their model and their numerical HMF “are larger at high $\nu$, where the best fit predicts more halos than we find in the simulations at $z=0$.” Since $\delta_{{\rm v}}(z)$ increases at low-$z$, whereas $\delta_{c}(z)$ decreases, it is clear that a HMF with the virialization threshold is more suitable. This is also represented by the better agreement between STv and E-HMFs at higher masses, which occurs for all redshifts and values of $\Omega_{m0}$ shown in figure 3.
One might ask [^1] We thank the referee for point this issue. whether the agreement between ST template (still assuming $\delta_{c}(z)$ as a threshold) and E-HMF can be improved if it is refitted in the same fashion as we did for the STv-HMF, i.e., assuming the E-HMF as a fiducial model and minimising the analogous $\chi^{2}$, given by equation (4.4). This refit indicates:
$a=0.7854\,,\,\,\,p=0.0990\,,\,\,\,A=0.3195\,.$ (4.7)
Note that the main difference with respect to the Despali et al. fit, equation (4.2), is lower $p$ value. The main effect of this new set of parameters is a better agreement between ST and E-HMFs at low masses. In figure 3, we refer to this new fit as STc. As can be seen, the behaviour at high masses is similar to the one for ST-HMF. When compared with this refit, the STv-HMF still performs better, especially at high masses. In particular, we can compare the values of $\chi^{2}$ for the best-fit parameters: STv-HMF has $\chi^{2}_{ST\rm{v}}=1472\times 10^{-3}$ and STc-HMF $\chi^{2}_{STc}=4325\times 10^{-3}$. Therefore we can conclude that lower $p$ values can improve the agreement between the ST-HMF template and E-HMF at low masses, both for $\delta_{c}(z)$ and $\delta_{\rm v}(z)$ thresholds. But at high masses $\delta_{\rm v}(z)$ is the main factor providing a better agreement between ST template and E-HMF.
It is important to note that large disagreements between any two distinct HMFs models is inevitable at the high-mass tail because any small difference between the parameters appearing in the exponential part will greatly amplify the differences at these mass scales. However, the actual consequences of this fact are usually not so drastic because high-mass objects are rare and demand very large simulations to be properly sampled, resulting in HMFs with decreasing accuracies with mass. Therefore, possible cosmological constraints and comparisons between HMFs have limited significance when considering the most massive halos at each epoch. The main message in the analysis portraited in figure 3 is the improved agreement between STv-HMF and the E-HMF with respect to the other versions of the ST template that use $\delta_{c}$. As we will further discuss in the next section, given this extended compatibility, the use of the STv-HMF is a good alternative to extrapolate the applicability of the E-HMF to models more complex than $\Lambda$CDM.
@@ 5 Beyond $\Lambda$CDM models
As a further consistency check of the use of $\delta_{{\rm v}}$ as the threshold for halo formation in the ST template, we will analyse the level of agreement between E, ST and STv-HMFs when computing the impact of smooth DE models with $w\neq-1$. HMFs for smooth DE with varying equation of state parameter have been studied by [7], where the effects were modelled by an explicit dependence on the matter growth rate, $f=d\ln\delta_{m}/d\ln a$. However, as shown in [8], this HMF has an overall worse agreement with the E-HMF than the ST-HMF. It is also discussed that the dependence on $f$ can be reparameterised in terms of $\Omega_{m}(z)$, which appears in equation (3.5), making use of growth index parametrization $f=\Omega_{m}^{\gamma}(z)$. Therefore, although the E-HMF has been fitted for the $\Lambda$CDM model, one can expect that the impact of $w\neq-1$ can be consistently captured by the explicit $\Omega_{m}(z)$ dependence [^2] In a private communication, Tiago Castro mentioned that an upcoming paper will show that the E-HMF can describe, within a few percent accuracy, the halo abundances in DE models with equation of state described by the $w0wa$ parametrization., and it will be used here as a basis of comparison.
@@ 5.1 Constant EoS
Figure 4: $R_{n}$, defined in (5.2), for STv, ST and E-HMFs (indicated in the plot legends) for dark energy models with $w_{0}=-0.8$ (left panel) and $w_{0}=-1.2$ (right panel) at $z=0$ (blue lines) and $z=0.5$ (orange lines). In the left panel, the disagreement between $R_{n}$ from STv and E-HMF is below $5\%$ for $M<2\times 10^{15}M_{\odot}h^{-1}$ at $z=0$ and $M<9.5\times 10^{14}M_{\odot}h^{-1}$ at $z=0.5$. In the right panel, the disagreement between $R_{n}$ from STv and E-HMF is below $5\%$ for $M<2\times 10^{15}M_{\odot}h^{-1}$ at $z=0$ and $M<9.5\times 10^{14}M_{\odot}h^{-1}$ at $z=0.5$.
An important feature to note in these examples is the clear impact of phantom ($w<-1$) and non-phantom EoS ($w>-1$) with respect to $\Lambda$CDM. For $w_{0}=-0.8$, the DE fluid has more energy density than $\Lambda$ at high-$z$, consequently $\Omega_{m}\left(z\right)$ is lower and the growth of matter fluctuations is relatively suppressed. As a consequence, we always have $R_{n}<1$. The opposite occurs for the phantom case of $w_{0}=-1.2$.
@@ 5.2 Time varying EoS
Although the impact of constant EoS is very clear, and all the three HMFs agree qualitatively about it, the situation for time-varying EoS can be more complex. As an example, we will analyse two pairs of $\left(w_{0},w_{a}\right)$ values, whose growth history changes in time with respect to the $\Lambda$CDM one. For $p_{1}=\left(w_{0}=-0.9,w_{a}=-0.3\right)$ the EoS is phantom at high-$z$ and non-phantom at low-$z$. For $p_{2}=\left(w_{0}=-1.104,w_{a}=0.3\right)$, the converse holds. It is interesting to note that the first kind of EoS evolution is supported by the recent DESI-BAO analysis [41].
Let us first analyse the relative linear growth history with respect to $\Lambda$CDM. In the left panel of figure 5, we show the ratio between the evolutions of $\sigma_{8}\left(z\right)$, given by
@@
For $p_{1}$, $\sigma_{8}(z)$ is higher than in $\Lambda$CDM at high-$z$. At lower $z$, the EoS becomes non-phantom and there is less growth with respect to $\Lambda$CDM. The opposite happens for $p_{2}$. In both cases, this relative transition in the growth behaviour occurs at $z_{t}\simeq 0.38$ (the value of $w_{0}$ in $p_{2}$ was adjusted to get approximately the same $z_{t}$ as for $p_{1}$).
Figure 5: Left panel: evolution of $R_{\sigma_{8}}$, defined in (5.4), for $p_{1}=\left(w_{0}=-0.9,w_{a}=-0.3\right)$ and $p_{2}=\left(w_{0}=-1.104,w_{a}=0.3\right)$. Right panel: evolution of $R_{c}$ and $R_{{\rm v}}$, defined in (5.5), for the same pairs of parameters.
Prior to $z_{t}\simeq 0.38$, the growth histories of the two examples are clearly enhanced or diminished with respect to $\Lambda$CDM. It is not expected that the nonlinear evolution encoded in the effective threshold of HMFs should present a distinct relative behaviour as seen for $R_{\sigma_{8}}$. Therefore, at $z>z_{t}$, the ratio between halo number densities, equation (5.2), should preserve the relations seen in the left panel of figure 5.
However, these relations between models based on the linear growth history is not always satisfied by HMFs. In figure 6 we show $R_{n}$ for the two models considered in this subsection. As can be seen, at $z=0.4$, slightly before $z_{t}\simeq 0.38$, only ST-HMF presents the expected behaviour: $R_{n}>1$ for $p_{1}$ and $R_{n}<1$ for $p_{2}$. This is easy to understand, given that the effective threshold in the ST-HMF depends on $\delta_{c}(z)$, which is nearly constant, the main impact in $R_{n}$ is due to $\sigma\left(M,z\right)$. For STv and E-HMF, the effective thresholds vary in time, but in a manner not totally consistent with the variation of $\sigma_{8}\left(z\right)$.
@@
Figure 6: Ratios of STv, ST and E-HMFs (indicated in the plot legends) for DE models with EoS parameters indicated at the plot titles at $z=0$ (blue lines) and $z=1$ (orange lines). In both panels, the disagreement between $R_{n}$ from STv and E-HMF barely exceeds $1\%$ for entire mass range shown, both at $z=0$ and $z=0.4$.
In order to clearly see that this undesirable effect is caused by the effective threshold, in the right panel of figure 5, we show the redshift evolution of the ratios
$R_{c}=\frac{D_{c\,\text{mod}}(z)}{D_{c\,\Lambda}(z)}\,\,\,\text{and }R_{{\rm v}}=\frac{\delta_{{\rm v\,}\text{mod}}(z)}{\delta_{{\rm v}\,\Lambda}(z)}\,,$ (5.5)
where $D_{c}$ is the time-dependent part of the effective threshold in the E-HMF, defined in (3.13). At high-$z$, both quantities are lower than unity for $p_{1}$ and higher than unity for $p_{2}$, as expected for enhanced and diminished growth histories with respect to $\Lambda$CDM, respectively. The inversion in the relative behaviour occurs much earlier than the one related to $\sigma_{8}\left(z\right)$, shown in the left panel of figure 5. After this inversion, $R_{\rm v}$ remains above or lower the unity, whereas $R_{c}\rightarrow 1$ as $z\rightarrow 0$. Moreover, $R_{c}$ reaches higher or lower values than $R_{\rm v}$, with a stronger time variation.
@@
Moreover, $R_{c}$ shows a second problem. Given that the redshift evolution of the effective threshold in the E-HMF mainly depends on $\Omega_{m}^{a_{z}}(z)$, for fixed $\Omega_{m0}$, $R_{c}$ for different models have the same value at $z=0$. This looks unnatural and is in disagreement with behaviour $R_{\sigma_{8}}$ shown in figure 5. Such behaviour would indicate that, although a specific model has more/less linear growth than $\Lambda$CDM at low-$z$, the effective threshold does not respond to this increase/decrease of growth at low-$z$. This issue might indicate that the parameter $a_{z}$ in the E-HMF should have some dependence on the EoS. On the other hand, $R_{{\rm v}}$ reproduces the behaviour of $R_{\sigma_{8}}$ at low-$z$, what looks more natural.
@@ 6 Conclusions
We have analysed the redshift dependence of the effective threshold in the E-HMF, proposed in [8], showing it is qualitatively equivalent to the virialization threshold, $\delta_{{\rm v}}(z)$, and incompatible with the evolution of the collapse threshold, $\delta_{c}(z)$, which has been used in many HMFs so far. This finding indicates that the abundance of halos in the Universe is better described by the physics of virialization, encoded in $\delta_{{\rm v}}(z)$, than physics of collapse, associated with $\delta_{c}(z)$.
We found that the redshift-dependent part of effective threshold in the E-HMF, $\Omega_{m}^{a_{z}}(z)\delta_{c}^{2}(z)$, can be accurately represented by a simple power-law function of the virialization threshold, given by $\alpha\delta_{{\rm v}}^{\beta}(z)$, which is also robust against variations of $\Omega_{m0}$, showing that the natural dependence of the threshold on $\delta_{{\rm v}}(z)$ is not a coincidence for a particular expansion history in the $\Lambda$CDM model. We found $\beta\simeq 4.37$, which shows that the effective threshold in the E-HMF grows faster than $\delta_{{\rm v}}^{2}(z)$ at low-$z$. We interpret this as an indication that a more realistic model for the nonlinear dynamics and virialization should provide faster growing $\delta_{{\rm v}}(z)$, in such a way that effective threshold is proportional to the square of this new function, as demanded by the analytical formulation of the HMF.
Exploring these findings, we have shown that a refitted ST-HMF with the substitution $\delta_{c}(z)\rightarrow\delta_{{\rm v}}(z)$ has better agreement with the E-HMF, which is also robust when varying $\Omega_{m0}$. Interestingly, this refit produce parameters $a$ and $p$ closer to $1$ and 00, respectively, making the STv-HMF more similar to the PS-HMF. Therefore, the STv-HMF is less dependent on ad-doc parameters, while providing better concordance with the E-HMF then the ST template with $\delta_{c}(z)$. In this sense, the STv-HMF preserves the universality features of the ST template with improved accuracy.
@@
Our main conclusion is that the virialization threshold is a sensible basic model for the description of the time-dependence of the effective threshold for smooth DE models. Given that $\delta_{{\rm v}}$ is computed by a physical model, in principle, it should provide a reliable manner to determine the impact models beyond smooth DE in halo abundances, such as clustering DE, or even for the scenarios of modified gravity and models beyond cold dark matter. Furthermore, at least in the case of clustering DE, it is not numerically possible to compute $\delta_{c}$ [29], thus the use of $\delta_{{\rm v}}$ is essential to explore the behaviour of such models with semi-analytic HMFs.